OpenStudy (anonymous):

(r^2 -5r- 2)(-r^2 +6r-6)

OpenStudy (anonymous):

Same as the last question.

OpenStudy (anonymous):

im confused by the r^2 and -r^2

OpenStudy (anonymous):

do they cancel

OpenStudy (anonymous):

You're multiplying not adding. They don't cancel.

OpenStudy (whpalmer4):

@brit3 \[r^2 = r*r\]\[-r^2 = -r*r\]so when you multiply those first terms together, you are forming \[(-r*r)*(r*r) = -r*r*r*r = -r^4\]